# Using Biot-Savart boundary conditions for unbounded external flow on Eulerian meshes

### v2

@@ 2.2 Fast evaluation of the Biot-Savart integral
$\mathcal{D}^{(4)}$
@@
Figure 2: Schematic of the multilevel approach to evaluate the Biot-Savart integral for a boundary point $\vec{x}$ over the nested domains $\mathcal{D}^{(1)}\cup\mathcal{D}^{(2)}\cup\mathcal{D}^{(3)}\cup\mathcal{D}^{(4)}$. The second domain half-width is shown, ${S}^{(2)}$. The computational domain $\Omega$ and its boundary $\partial\Omega$ and the immersed body $\mathcal{B}$ are shown. We also show a schematic of the coarsening of the grid between the domains. Domain and cell size are only for representation proposes.
The aim of this novel boundary conditions is to reduce the computational cost compared to a large grid by providing an accrate estimate of the far field conditions. A naive evaluation of Eq. 8 would require $O(N)$ operations for each boundary point, meaning $O(N^{1+\cal S})$ cost overall, potentially overwhelming any savings from reducing the size of the domain. In vortex particle-mesh methods, the integral is usually evaluated with Fast Multipole Method that allows reducing this cost to $O(\log N)$ operations. Here, we use a multilevel algorithm to achieve the same speed-up in our Eulerian setting.
We construct the multilevel Biot-Savart operator for each point $\vec{x}$ on the domain boundary $\partial\Omega$ by representing the complete domain $\Omega$ with a combination of gradually coarser meshes
@@
where level $1$ is the finest domain and $l$ is the coarsest, and each domain is non-overlapping and their union fully covers $\Omega$, Fig 2. In this work, we use nested rectangular boxes of half-width $S^{(i)}$ centered on $\vec{x}$, and scale this width with the cell size which doubles at each level, i.e. $h^{(i)}=h^{(1)}2^{i-1}$. Therefore the domains are determined by $\vec{x}$ and a non-dimensional half-width
$\tilde{S}=\frac{{S}^{(l)}}{2^{l-1}},$ (11)
which is equal on every level. This domain decomposition is chosen so that the contributions to Eq. (8) are more accurately calculated when close to $\vec{x}$, and the smaller contributions further from $\vec{x}$ are quickly estimated on the coarser levels.
With this set of $l$ non-overlapping domains $\mathcal{D}^{(i)}$, the integral in Eq. 8 can be replaced with a sum over all the contribution from the domains
$f(\vec{x};\vec{\omega},\Omega)\approx\sum_{i=1}^{l}f(\vec{x};\vec{\omega}^{(i)},\mathcal{D}^{(i)}).$ (12)
Once the multilevel field $\omega^{(i)}$ is computed, the total number of operations to evaluate Eq. 12 for a simulation of dimension $d=2$ or $3$ is therefore $O(l\tilde{S}^{d})=O(\log N)$ since $\tilde{S}$ is constant and $l=O(\frac{1}{d}\log N)$.
In practice, we use the circulation over a cell face $\Gamma=\omega h^{2}$ instead of the vorticity in the Biot-Savart formulation. The circulation on a coarser level is calculated by uniformly pooling the circulation on the level above
$\Gamma^{(i+1)}=\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (13)
where $\Gamma^{(1)}$ is the original finest level field and the pooling $\mathcal{P}$ for a cell at level $i$ simply sums over the corresponding subcells in level $i-1$, see [26]. Thus, our multi-level field $\Gamma^{(i)}$ conserves the total circulation to machine precision, as recommended by [2] and others.
Finally, we derive the error induced by this multilevel approach, starting with the 2D-case and extending the result to 3D. The velocity at location $\vec{x}$ induced by the circulation from a single cell on level $i+1$ is given by (see Fig. 2)
After rearrangement, the first term equals Eq. (14). The second term is the error due to the multilevel approximation. Its magnitude is
2.3 Biot-Savart & Projection Algorithm
The momentum update using the new Biot-Savart domain condition proceeds as before: forming the intermediate velocity with Eq. 2 and projecting it to be divergence free with Eq. 6. However, the new boundary condition introduces a coupling between the body and domain boundaries. To see this, we write the equation for the boundary velocity on $\vec{x}_{b}\in\partial\Omega(N^{\cal S})$ at the new time-step as
where we’ve substituted Eq. 6. The last term in Eq. 20 is non-zero on cells touching the body boundary $\partial\cal B$ because, while the gradient of a field cannot induce curl in an unbounded fluid region (×$\nabla\times\nabla\phi\equiv 0$), the pressure boundary condition on the immersed body generates a thin layer of vorticity on $\partial\cal B$. This contribution is unknown until after the Poisson system of Eq. 5 is solved, but updating the normal velocity on $\partial\Omega(N^{\cal S})$ will change the source term for Eq. 5 on cells touching the domain boundary, requiring a new pressure solution.
This coupling between the domain and body boundaries results in a mildly non-linear fixed-point system. This weak non-linearity can be solved efficiently with a deferred correction approach, which explicitly updates the source term of the Poisson problem during pressure solver iteration. We use a Multi-Grid solver for the Poisson system, as described in [26], and updating the domain velocities and corresponding Poisson residuals at the end of each V-cycle was sufficient to ensure the final pressure and velocity field satisfy both the momentum and divergence-free conditions without increasing the number of solver iterations. Note that this coupling is only apparent in immersed boundary methods if the pressure boundary conditions are correctly imposed on the immersed body during the projection step; see for example [27, 20]. This approach differs from the one in [10] for unbounded Poisson problems with embedded Dirichlet boundaries as we directly adjust the residual of the (iterative) Poisson solver and we do not iterate over the coupled Poisson/interior boundary problem. It is also different to the James-Lackner algorithm for the same reason [13].
Lastly, we note that the small multilevel error discussed in the previous section means that global mass conservation is no longer guaranteed to machine precision on the domain boundary. As such, we apply a uniform correction to the velocity computed with Eq. 20
$\vec{u}(\vec{x}_{b})=\vec{u}(\vec{x}_{b})-\frac{\oint_{\partial\Omega}\hat{n}\cdot\vec{u}\text{ d}S}{\oint_{\partial\Omega}\text{ d}S}.$ (21)
ensuring a globally divergence-free field is passed to the pressure projection method on each iteration.
3 Multilevel reconstruction validation
We validate the multilevel reconstruction algorithm presented above using 2D and 3D analytic vorticity fields. We also study the multilevel reconstruction error and the speedup generated by selecting the correct kernel radius from the above analysis.
Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^2] https://github.com/weymouth/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [22] that can execute both on CPU and GPU [28]. To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory). The complete code and instructions needed to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
3.1 2D validation: Lamb Chaplygin vortex-dipole
The velocity field generated by the Lamb-Chaplygin dipole is defined by the scalar stream function
$10^{-1}$
$10^{-2}$
$10^{-3}$
$10^{-4}$
$10^{-5}$
$10^{-6}$
$a)$ $S=2$
$c)$ $S=8$
$d)$ $S=16$
$f)$ $S=128$
$b)$ $S=4$
$e)$ $S=32$
Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size.
We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result. We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$. The dependency of the absolute error in the reconstructed velocity field on the radius of the multilevel refinement criterion are shown in Fig. 3. Setting the refine size $S$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$. This error depends primarily on the distance from the vortex, Fig. 3(f). Progressively reducing the refinement size $S$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
This is further quantified in Fig. 4(a), which shows the maximum error in the reconstructed velocity field at a distance $d$ from the vortex from the Lamb-Chaplygin vortex. The size $S$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $S\leq 2^{3}=8$. Fig. 4(b) reports the measured computational speed-up achieved by the multilevel method relative to setting $S=\Omega$, demonstrating up to 300$\times$ acceleration.
3.2 3D validation: Hill vortex
and the stream function
with the corresponding errors and speed-up shown in Figure 5. Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth 2D Lamb-dipole.
As argued in Sec. 2.3 the multilevel error decreases extremely quickly as $S$ is increased for three-dimensional flows. This is due to the scaling of the Biot-Savart kernel itself. The 3D case demonstrates a stronger sensitivity to the distance to vortex, as expected from the error estimate. Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a max observed speed-up of nearly 20,000$\times$ using the multilevel method.
@@ 4 Confined vorticity applications
Next, we validate the application of the Biot-Savart conditions to a set of impulsive flow problems. The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
@@
The flow around a 2D or 3D disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$. As a result, an analytical expression for the added mass coefficient can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
@@
We perform 2D and 3D simulations of the flow around a plate/circular disk immersed in a domain of dimension $1.5D\times 1.5D(\times 1.5D)$. From potential flow theory, the added-mass coefficient for an impulsively started flow is
@@
We note here that the 2D and 3D problems possess axes of symmetry that would allow only a portion of the domain to be modeled. This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each vortex cell, see A. However, to avoid unnecessary complications in this manuscript, we model the full plate and do not use the problem’s symmetries.
@@
Figure 6 shows the vorticity and pressure field at three consecutive times during the disk’s motion on the midplane of the 3D circular disk. The top row shows the results with the standard reflective boundary conditions (the entire computational domain is shown), and the bottom row shows the flow obtained with the Biot-Savart boundary conditions. High blockage effects are apparent for the reflective boundary conditions, which generate a stronger shear layer that feeds a stronger vortex located further downstream than the Biot-Savart results. The error in the pressure field due to the homogeneous Neumann condition placed close to the immersed body generates pressure peaks twice as large as with the Biot-Savart boundary conditions.
@@
Figure 6: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions (top) and the Biot-Savart boundary condition (bottom). A slice through the middle of the entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
@@
Fig. 7 shows the time trace of the pressure force acting on the accelerated disk, as well as the equivalent 2D simulation on an accelerating flat plate for comparison. The difference in the pressure field is found again in the difference in the forces. Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it. For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces for all $t$.
@@
Figure 7: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6
@@
We qualitatively compare the shape of the vortex ring forming behind our accelerating square plate with the flow visualization of the same cases in [29], who used dye coloring to visualize the wake forming behind the accelerating plate. Results are presented in Figure 9-8 for four instants during and after the acceleration phase.
@@
5 Vortex-street applications
Next, we study the accuracy of the new Biot-Savart conditions on flow cases with vortex wakes that are truncated by the domain boundary, introducing an additional modelling error. We use two classical examples of vortex-street flows: (1) the flow around a 2D cylinder at a Reynolds number $\text{Re}=200$, which results in the classical van Kàrmàn wake, and (2) the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
5.1 Flow around a 2D cylinder at $\text{Re}=200$
We start our analysis with the flow behind a 2D circular cylinder of diameter $D$. The cylinder is immersed in a viscous fluid with free-stream velocity $U=(1,0)$. The viscosity is set such that the Reynolds number is $\text{Re}=\frac{UD}{\nu}=200$. We perform a domain study by varying the domain sizes from $5D\times 3D$ up to $30D\times 24D$ and, for each domain, compare the reflective and Biot-Savart boundary conditions.
Figure 10 shows the total drag force acting on the cylinder for the various domain sizes and boundary conditions. The Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forcesregardless of the domain size, including the tiny $5Dx3D$ domain which captures no more than two wake vortices at any time. Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution with a domain 48x smaller.
Additionally, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7D\times 7D$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions. Results are also presented in Fig. 10 and are almost indistinguishable from the results using the aligned domain.
In addition to the total drag forces acting on the cylinder, we look at the vortex shedding frequency in the wake. Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$. The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, adding a non-physical domain-size dependance on the shedding frequency.
Figure 10: (left) Drag force acting on the circular cylinder at $\text{Re}=200$ for various combinations of boundary conditions and domain size. The rotated domain consists of a $4D\times 4D$ domain, with the cylinder centered at $1.5D\times 1.5D$ and a free stream velocity of $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$. (right) The frequency spectrum of the wake of the cylinder for the same combination of boundary conditions and domain size shows the classical shedding frequency at $St=0.3$.

### v3

@@ 2.2 Fast evaluation of the Biot-Savart integral
The aim of Biot-Savart boundary conditions is to reduce the computational cost compared to a large grid by providing an accurate estimate of external flows using small domains. While the boundary conditions are agnostic of the method used to compute the integral in Eq. 8, a naive convolution over $N$ vorticity “source” locations $\vec{y}$ would require $O(N)$ operations for every “target” velocity $\vec{u}(\vec{x})$ on the boundary. The overall $O(N^{1+\cal S})$ cost would likely overwhelm any benefit of the reduced grid size, motivating a faster Biot-Savart evaluation approach.
Highly precise Fast Multipole Methods (FMMs) have been developed to a reduce this cost to $O(\log N)$ or even $O(1)$ operations per target, [24, 25, 26]. Our context allows for significant specialization and simplification to a standard FMM. (i) A general FMM considers a general set of source locations, and they are typically optimized to evaluate on the same set of targets. In contrast, our sources are equally spaced over the domain and our targets are equally spaced on it’s boundary, meaning we have only $O(N^{\cal S})$ targets which are exclusively exterior to the source set. (ii) The high error tolerance discussed in the section above means truncating the multipole has no effect on the satisfaction of the divergence-free constraint (set by the Poisson solver tolerance) and little effect on the solution accuracy overall.
@@
$\mathcal{P}^{(3)\to(4)}$
$\mathcal{D}^{(2)}$
$\mathcal{B}$
Figure 2: Schematic of the multilevel approach to evaluate the Biot-Savart integral for a boundary point $\vec{x}$ over the nested subdomains $\mathcal{D}^{(1)}\cup\mathcal{D}^{(2)}\cup\mathcal{D}^{(3)}\cup\mathcal{D}^{(4)}$. The second domain half-width is shown, ${S}^{(2)}$. The computational domain $\Omega$ and its boundary $\partial\Omega$ and the immersed body $\mathcal{B}$ are shown. We also show a schematic of the coarsening of the grid between the domains and the multilevel pooling $\mathcal{P}^{(3)\to(4)}$ of the cell’s circulation (circular-colored markers).
Based on this context, we use a simple multi-level vorticity source clustering method which leverages their uniform distribution in our Eulerian domain. We construct a multi-level vorticity field $\omega^{(i)}$ where level $i=1$ is the finest and $l$ is the coarsest and the cell size doubles with each level $h^{(i+1)}=2h^{(i)}$. The levels are filled from finest to coarsest recursively by uniformly pooling the circulation $\Gamma=\omega h^{2}$ on the level above
$\Gamma^{(i+1)}=\mathcal{P}^{(i)\to(i+1)}\Gamma^{(i)}$ (17)
where the pooling $\mathcal{P}$ for a cell at level $i+1$ simply sums over the corresponding sub-cells in level $i$, see Fig. 2. Thus, our multi-level field conserves the total circulation to machine precision, as recommended by [2] and others. Filling the entire multi-level source field requires $O(N)$ operations.
In this work, we present results from two boundary velocity reconstruction approaches using the multi-level source clustering described above:
  1. 1.
    tree-sum: In the first approach, the integral over the original domain $\Omega$ in Eq. 8 is replaced with a sum over the contribution from a sequence of subdomains
@@
    where we define the sub-domains as nested rectangular boxes of half-width $S^{(i)}=h^{(i)}\tilde{S}$ centered on the target location $\vec{x}$, Fig 2. As the sub-domains are non-overlapping and fully cover $\Omega$, Eq. 8 becomes
    $\vec{u}(\vec{x})\approx\sum_{i=1}^{l}f(\vec{x};\vec{\omega}^{(i)},\mathcal{D}^{(i)}(\vec{x})).$ (19)
    This corresponds to a direct sum for each target $\vec{x}$ over an oct-tree source field in 3D (or a quad-tree field in 2D). Therefore, we call this the “tree-sum” approach. As each subdomain has $O(\tilde{S}^{n})$ sources and the number of levels $l=O(\log N)$, the tree-sum requires $O(\tilde{S}^{n}\log N)$ operations per target.
  2. 2.
    $\ell$M: In the second approach, we also cluster the domain boundary target points $\vec{x}^{\,(i)}$, doubling the spacing at each level. We then define the level-$i$ kernel interaction velocities as
    $\vec{v}^{\,(i)}(\vec{x}^{\,(i)})=f(\vec{x}^{\,(i)};\vec{\omega}^{(i)},\mathcal{D}^{(i)}(\vec{x}^{\,(i)})).$ (20)
    The target velocities are then unpooled recursively from the coarsest level to the finest
    $\vec{u}^{\,(i)}\approx\vec{v}^{\,(i)}+\mathcal{P}^{(i+1)\to(i)}\vec{u}^{\,(i+1)}$ (21)
    where $u^{\,(l+1)}=0$ and the unpooling operator is bi-linear interpolation to the target locations $\vec{x}^{\,(i)}$. As the interaction velocities are independent at each level, they require a constant $O(\tilde{S}^{n})$ operations per target. Therefore, we call this the Fast Multi-level Method ($\ell$M).
A key difference between these approaches and a classic FMM is that the contributions from each clustered source point are obtained by evaluating Eq. 9 directly and not through a classical or kernel-independent multipole expansion, as done in [25, 12]. As such, the domain half-width $\tilde{S}$ controls the reconstruction error and A proves that this error is bounded and inversely proportional to the number of subdomain points $\tilde{S}^{n}$. The results sections below demonstrate that these approaches are more than sufficient for high-accuracy external flow predictions.
Finally, we note that we can parallelize the velocity reconstruction to further accelerate the method since each Biot-Savart kernel evaluation is independent. Our implementation applies GPU or CPU multi-threading to parallelize the calculation and pooling of the vorticity across the sources, as well as the kernel function evaluation and velocity unpooling across the targets.
@@ 3 Confined vorticity applications
We first validate the application of the Biot-Savart boundary conditions using a set of impulsive flow problems. The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^1] https://github.com/WaterLily-jl/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [27] that can execute both on CPU and GPU [28]. To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory). The complete code and instructions needed to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
3.1 Flow around a 2D cylinder at $\text{Re}=550$
We start our analysis with the flow behind a 2D circular cylinder of diameter $D$. The cylinder is immersed in a viscous fluid and impulsively accelerated to free-stream velocity $U$. The viscosity is set such that the Reynolds number is $\text{Re}=\frac{UD}{\nu}=550$. We set $D=128$ cells for all tests and use a $2\times 1$ rectangular domain, varying the width $W$ from $=2D\ldots 24D$. Figure 3(left) shows the vorticity field on the smallest domain at convective time $t^{*}=tU/D=4$.
Figure 3(right) compares the predicted drag coefficient using the new Biot-Savart boundary conditions to the analytical solution [29] (for $t^{*}\ll 1$) and two sets of reference data from high-resolution vortex-based methods [30, 31]. The new condition is able to accurately predict the total drag force on the cylinder, matching the theoretical and vortex-based methods even when using the smallest $4D\times 2D$ domain with 50% blockage. Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the external flow solution. Even using a domain 100 times larger produces noticeable deviations for $t^{*}>3.5$.
Figure 3: Impulsively started circular cylinder at $\text{Re}=550$ validation case. (Right) Vorticity field $\omega D/U$ at convective time $tU/D=4$ over the full $2D\times 4D$ domain. (left) Early time history of the drag coefficient $C_{d}=F_{d}/\frac{1}{2}\rho U^{2}D$ for various domain blockage ratios $D/W$. The present Biot-Savart BCs match the theoretical and vortex-based methods even with 50% blockage.
@@ 3.1 Flow around a 2D cylinder at $\text{Re}=550$
The flow around a 2D or 3D disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$. As a result, an analytical expression for the added mass coefficient can be used to further validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
@@
We perform 2D and 3D simulations of the flow around a plate/circular disk immersed in a domain of dimension $1.5D\times 1.5D\,\,(\times 1.5D)$ with a resolution of 86 cells per diameter ($D$). From potential flow theory, the added-mass coefficient for an impulsively started flow is
@@
We note here that the 2D and 3D problems possess axes of symmetry that would allow only a portion of the domain to be modeled. This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each vortex cell, see B. However, to avoid unnecessary complications in this manuscript, we model the full plate and do not use the problem’s symmetries.
@@
Figure 4 shows the vorticity and pressure field for three consecutive times during the disk’s motion on the midplane of the 3D circular disk. The top row shows the results with the standard reflective boundary conditions (the entire computational domain is shown), and the bottom row shows the flow obtained with the Biot-Savart boundary conditions. High blockage effects are apparent for the reflective boundary conditions, which generate a stronger shear layer that feeds a stronger vortex located further downstream than the Biot-Savart results. The error in the pressure field due to the homogeneous Neumann condition placed close to the immersed body generates pressure peaks twice as large as with the Biot-Savart boundary conditions.
@@
Figure 4: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions (top) and the Biot-Savart boundary condition (bottom). A slice through the middle of the entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega R/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
@@
Fig. 5 shows the time trace of the pressure force acting on the accelerated disk, as well as the equivalent 2D simulation on an accelerating flat plate for comparison. The impact of the error in in the pressure field is apparent in the difference in the forces. Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it. For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces for all $t$.
@@
Figure 5: Instantaneous pressure forces normalized by added mass acting on the 2 and 3D disk accelerated from rest for (a) the two different boundary conditions and (b) for different Biot-Savart kernels (tree-sum and $\ell$M) and different subdomain half-width $\tilde{S}$ for the 3D disk. The convective time corresponds to the snapshots shown in Fig. 4.
Next we demonstrate the low sensitivity of the Eulerian flow solver using the simple partitioned approach of Eq. 16 to the implementation details of the Biot-Savart kernel summation. Fig. 5(b) compares the instantaneous drag force acting on the 3D disk using the tree-sum and $\ell$M approach over a range of clustering subdomain sizes. The results are extremely consistent, with any $\tilde{S}\geq 2$ predicting identical forces and only the smallest possible half-width $\tilde{S}=1$ leading to a few percent increase in the predicted peak force at $t^{*}=2$. In addition, the forces are identical using the two approaches despite the additional unpooling step of the $\ell$M.
This minimal sensitivity means that we can maintain accuracy while maximizing the reconstruction speed. For the remainder of the manuscript we will use the $\ell$M approach with $\tilde{S}=2$ for 3D flows, and $\tilde{S}=4$ for 2D. We therefore use $O(N)$ operations for clustering the sources and $O(N^{\cal S})$ operations for reconstructing the target velocities for each V-cycle of the pressure solver, which itself is $O(N)$. In practice, we find that using Biot-Savart boundary conditions on small domains in 2D can double the simulation speed compared to using reflection condition on the same domain because the artificially fast flow speeds shown in Fig. 4 decrease the possible time-step. In 3D we find simulation times are basically identical using Biot-Savart boundary conditions or reflection conditions on the same domain, with the small reconstruction cost and increased time-steps roughly canceling.
@@
We qualitatively compare the shape of the vortex ring forming behind our accelerating square plate with the flow visualization of the same cases in [32], who used dye coloring to visualize the wake forming behind the accelerating plate. Results are presented in Figure 6, 7 for four instants during and after the acceleration phase.
@@
4 Developed vortex-wake applications
Next, we study the accuracy of the new Biot-Savart conditions on flow cases with vortex wakes that are truncated by the domain boundary, introducing an additional modeling error. We use two canonical developed vortex-wake examples: (1) the flow around a sphere at $\text{Re}=3700$, and (2) the reversed van Kàrmàn (or propulsive) wake generated by a heaving 2D airfoil over a range of Strouhal numbers.
4.1 3D flow around a sphere
Figure 8: Sphere flow vortex wake and drag coefficient at $Re=UR/\nu=3700$ with a resolution of 44 cells per radius. (top) Wake visualized by a volumetric rendering of $|\omega|R/U$ at time $tU/R=53$ with the full $3.6D\times 1.5D\times 1.5D$ domain shown. Instantaneous (right) and time-averaged (left) drag coefficients for three domains, where $A$ is the domain frontal area. The time-averages drag is obtained by averaging the last 100 convective times.
The flow around a sphere is used to validate the coupled momentum-Biot-Savart update for developed 3D flows. A sphere of radius $R$ is immersed in a viscous incompressible flow. The Reynolds number is set to $Re=UR/\nu=3700$, we use a resolution of 44 cells per radius and use a domain size of $5C/2\times C\times C$ cells varied over $C\in[128,192,256]$. The flow is impulsively started from rest to a uniform velocity $U$.
We perform the simulations until convective time $tU/R=200$ and Figure 8 shows the drag coefficient when using the Biot-Savart boundary condition is insensitive to the domain size - even for domains only 50% wider than the sphere diameter. The averaged drag over the last 100 convective time for all blockage ratios compares favorably to the DNS and LES simulation results obtained in [33, 34].